I need to know what the answer to this equation is?
x^-5
-----=
y^-6

Answer by richwmiller(17219)

A negative exponent just means that the base is on the wrong side of the fraction line, so you need to flip the base to the other side. For instance, "x–2" (ecks to the minus two) just means "x2, but underneath, as in 1/(x2)".
* Write x–4 using only positive exponents.
x^(-4) = 1/x^4


what does negative exponent mean?
x^(-3) / x^(-7) = x^7 / x^3
x^-5
-----=
y^-6

y^6/x^5